OpenStudy (anonymous):

Matt is showing his work in simplifying (6.2 − 1.6) − 4.4 + 7.8. Identify any error in his work or reasoning. Step 1: (6.2 − 1.6) − 4.4 + 7.8 Step 2: 6.2 + (− 1.6 − 4.4) + 7.8 (commutative property) Step 3: 6.2 − 6 + 7.8 Step 4: −14 − 6 = 8 (commutative property) Step 5: 14 − 6 = 8 He wrote commutative instead of distributive in Step 4. He wrote commutative instead of distributive in Step 2. He wrote commutative instead of associative in Step 4. He wrote commutative instead of associative in Step 2.

OpenStudy (anonymous):

@JoannaBlackwelder

OpenStudy (joannablackwelder):

Do you know what the commutative property is?

OpenStudy (joannablackwelder):

@realmadrid7 ?

OpenStudy (anonymous):

a + b = b + a

OpenStudy (joannablackwelder):

Good! Does step 2 use that property?

OpenStudy (anonymous):

yes

OpenStudy (joannablackwelder):

It looks like the numbers are in the same order to me.

OpenStudy (joannablackwelder):

All that changed were the parentheses, right?

OpenStudy (anonymous):

yea

OpenStudy (joannablackwelder):

No negatives changed. One subtraction got changed to adding a negative, but that is not the commutative property.

OpenStudy (joannablackwelder):

So, what is the property that says you can move the parentheses?

OpenStudy (joannablackwelder):

(a+b)+c=a+(b+c) Which property is this?

OpenStudy (anonymous):

distrbutive property right ?

OpenStudy (joannablackwelder):

No, that is a(b+c)=ab+ac

OpenStudy (joannablackwelder):

?

OpenStudy (anonymous):

assosative property

OpenStudy (joannablackwelder):

Right :)

OpenStudy (anonymous):

mybad lol

OpenStudy (joannablackwelder):

haha, no worries

OpenStudy (anonymous):

so it would be d right

OpenStudy (anonymous):

@realmadrid7

OpenStudy (anonymous):

Do you remember if d was right?